A Bibliography of Publications of Shachar Fleishman

Shachar Fleishman School of Computer Science Tel-Aviv University Schriber Building, Room 002 Tel Aviv, 69978 Israel

Tel: +972-3-6405360 (University), +972-3-5732940 (home) FAX: n/a

E-mail: shacharf@tau.ac.il (Internet)

27 November 2011 Version 1.05

Abstract

This bibliography records publications of Shachar Fleishman.

Title word cross-reference

Bilateral [FDCO03].

California [EJV01]. Computing [ABCO^{+}03].

Denoising [FDCO03]. Diego [EJV01].

Mesh [FDCO03].

October [EJV01].

Point [ABCO^{+}01]. proceedings [EJV01]. Progressive [FACOS03].

Rendering [ABCO^{+}03].

San [EJV01]. Set [ABCO^{+}01]. Surfaces [ABCO^{+}01].

References

Alexa:2001:PSS

[ABCO^{+}01] M. Alexa, J. Behr, D. Cohen-Or, S. Fleishman, D. Levin, and C. T. Silva. Point set surfaces. In Ertl et al. [EJV01], pages 21– 28. ISBN 0-7803-7200-X. URL

                                                   http://ieeexplore.ieee.org/
                                                   servlet/opac?punumber=7631.

Alexa:2003:CRP

[ABCO^{+}03] M. Alexa, J. Behr, D. Cohen-Or, S. Fleishman, D. Levin, and C. T. Silva. Computing and rendering point set surfaces.

IEEE Transactions on Visualization and Computer Graphics, 9 (1):3–15, January 2003. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ertl:2001:PSS

[EJV01] Thomas Ertl, Ken Joy, and Amitabh Varshney, editors. Visualization 2001: proceedings: October 21–26, 2001, San Diego, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, October 2001. ISBN 0-7803-7200-X. URL

http://ieeexplore.ieee.org/
servlet/opac?punumber=7631.

Fleishman:2003:PPS

[FACOS03] Shachar Fleishman, Marc Alexa, Daniel Cohen-Or, and Claudio T. Silva. Progressive point set surfaces. ACM Transactions on Graphics, 22(4), 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fleishman:2003:BMD

[FDCO03] Shachar Fleishman, Iddo Drori, and Daniel Cohen-Or. Bilateral mesh denoising. ACM Transactions on Graphics, 22(3):950– 953, 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).